OpenStudy (lgbasallote):

Show that if n is an integer and n^3 + 5 is odd then n is even

OpenStudy (lgbasallote):

if i use contraposition, then that means i should assume n is odd first

OpenStudy (lgbasallote):

and n^3 + 5 is even

OpenStudy (lgbasallote):

so if n is odd... n = 2k + 1

OpenStudy (lgbasallote):

then if i substitute (2k+1)^3 + 5

OpenStudy (lgbasallote):

i assume this to be even so (2k + 1)^3 + 5 = 2n

OpenStudy (anonymous):

If n^3+5 is odd then n^3 is even thus n is even.

OpenStudy (lgbasallote):

that doesn'r make sense....

OpenStudy (anonymous):

Why?

OpenStudy (lgbasallote):

1. how can you concur n^3 is even from n^3 + 5 2. im using contraposition

OpenStudy (anonymous):

Sum of two odd numbers is always even......So 5 is odd and n^3+5 is also odd.Thus, n^3 must be even.

OpenStudy (lgbasallote):

but n^3 + 5 is the sum....not n^3

OpenStudy (anonymous):

Ya...... n^3+5 is the sum which is odd AND 5 is also odd....... which means n^3 must be even.

OpenStudy (lgbasallote):

you said sum of two odds is even

OpenStudy (anonymous):

Assume n odd -> n^3 odd

OpenStudy (lgbasallote):

n^3 is even 5 is odd how does that make sum of two odds

OpenStudy (anonymous):

Actually "Assume n odd -> n^3 odd" QED.

OpenStudy (anonymous):

Because sum of odd and even is always odd.

OpenStudy (lgbasallote):

better. however, like i said im using contraposition

OpenStudy (lgbasallote):

but i'll have to prove n^3 is odd when n is odd first...

OpenStudy (anonymous):

Why use contraposition?

OpenStudy (lgbasallote):

it was the specifics of the problem

OpenStudy (lgbasallote):

and also to teach myself versatility

OpenStudy (anonymous):

What kind of method is that......never heard of it.

OpenStudy (lgbasallote):

assume p and q are false therefore \(\neg q \rightarrow \neg p \equiv T\)

ganeshie8 (ganeshie8):

if a, then b => if not b, then not a

OpenStudy (lgbasallote):

so i suppose my next step is to prove n^3 is odd

OpenStudy (anonymous):

Wants to show If n not even, then blah is odd Going all round the houses.....

OpenStudy (anonymous):

n^3+5=odd and we know if we add two odd num. so we get even so n^3 always be even so n=even

OpenStudy (lgbasallote):

wait no...that's not the step

OpenStudy (lgbasallote):

@vipul92 contraposition

OpenStudy (anonymous):

I think that was what I did.

ganeshie8 (ganeshie8):

that would do : if n is odd, then n^3 is odd

OpenStudy (anonymous):

@vipul92

OpenStudy (lgbasallote):

1. how to prove n^3 is odd 2. the aim is to prove n^3 + 5 is even

OpenStudy (lgbasallote):

so i had (2k + 1)^3 + 5 = 2n 8k^3 + 12k^2 + 6k + 1 + 5 = 2n 8k^3 + 12k^2 + 6k + 6 = 2n 2(4k^3 + 6k^2 + 3k + 3) = 2n let 4k^3 + 6k^2 + 3k + 3 = x 2x = 2n QED

OpenStudy (mayankdevnani):

@igbasallote go through this link

OpenStudy (anonymous):

as we know the cube of any odd is odd and for any even is even

OpenStudy (lgbasallote):

@mayankdevnani i'd rather not have spoilers

OpenStudy (mayankdevnani):

what

OpenStudy (lgbasallote):

i don't want to see the future steps. it spoils the fun

OpenStudy (anonymous):

I think she wants to work it though....

OpenStudy (lgbasallote):

she?

OpenStudy (anonymous):

He, whatever...

OpenStudy (mayankdevnani):

ok

OpenStudy (lgbasallote):

just because i hate math doesn't make me a girl...just saying....

OpenStudy (anonymous):

Proof?

OpenStudy (lgbasallote):

anyway...did i do the contraposition right?

OpenStudy (anonymous):

Yes, u did (still think that's the long way round, though)

OpenStudy (lgbasallote):

what about contradiction? i just assume one proposition is false right?

OpenStudy (anonymous):

It's immediate by assuming n odd, then n^3 is odd...

OpenStudy (lgbasallote):

im not as good as you in math

OpenStudy (anonymous):

For the reason along the lines you used in the other proof..

OpenStudy (lgbasallote):

anyway...proving by contradiction would be... assume n = odd then i prove n^3 + 5 is odd right?

OpenStudy (anonymous):

Well u know odd by odd is odd, right (usually this can be assumed) so If n is odd so is n^3

OpenStudy (lgbasallote):

true...but as you may have noticed....i prefer going step by step...don't worry, you'll see me skipping a lot of steps when i'm sure i know proving already

OpenStudy (anonymous):

And your question already gives n^3 +5 as odd

OpenStudy (lgbasallote):

anyway...is my assumption on proving by contradiction right?

OpenStudy (anonymous):

No, because of what I just said,,,

OpenStudy (lgbasallote):

n^3 + 5 is odd?/

OpenStudy (anonymous):

The assumption here is that n is odd and we want to see if this assumption leads to contradiction.

OpenStudy (lgbasallote):

yes...

OpenStudy (anonymous):

So the first conclusion from this assumption is that n^3 is odd.

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

And so n^3 +5 is even (odd + odd is even), which is a contradiction.

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

although i just asked if my assumption for contradiction was right...

OpenStudy (anonymous):

So n is not odd, it is even QED.

OpenStudy (anonymous):

although i just asked if my assumption for contradiction was right... and I explained to you that the assumption is simply that n is odd

OpenStudy (anonymous):

Problem- show n even Assume - n odd Find- a contradiction Therefore - assumption is wrong

OpenStudy (lgbasallote):

so my assumption was right.

OpenStudy (lgbasallote):

by "assumption was right" i meant that it'sthe assumption i use for contradiction

OpenStudy (anonymous):

U said "assume n = odd then i prove n^3 + 5 is odd" This is incorrect.

OpenStudy (lgbasallote):

well i am going to prove by contradiction..

OpenStudy (lgbasallote):

so in a way my goal is to prove that n^3 + 5 is odd and contradict it

OpenStudy (anonymous):

You cannot prove n^3 +5 is odd because n^3 +5 is given as odd already.

OpenStudy (lgbasallote):

you dodn't have to take every word literally

OpenStudy (anonymous):

I'm not sure what you want me to say, the assumption (ie something not already given) u make is that n is odd and go from there....

OpenStudy (lgbasallote):

although i admit my wording is poor. i was just asking about the sense...

OpenStudy (lgbasallote):

yes that's what i was asking

OpenStudy (lgbasallote):

i was just asking whether i just assume one to be wrong

OpenStudy (anonymous):

You shouldn't make any prior judgement about where your assumption will lead....

OpenStudy (lgbasallote):

i didn't make any prior judgements..

OpenStudy (anonymous):

You simply hope that it will lead to something useful for your argument.

OpenStudy (anonymous):

If your are looking for the usual wording then it would be something like: "For the purpose of deriving a contradiction, we assume that n is odd...."

OpenStudy (lgbasallote):

that seems too formal and long...and like you've said before, you hate long methods when it can be done simply

OpenStudy (anonymous):

True, it depends on the situation (is it a test, what can u assume, what do you have to prove, etc)

OpenStudy (anonymous):

At some point, you leave behind the basic number properties and just concentrate on the logic and reasoning.

OpenStudy (anonymous):

Now your current question is iff so both directions...

OpenStudy (lgbasallote):

i believe in displacement over distance